# Joint Hierarchical Representation Learning of Samples and Features via Informed Tree-Wasserstein Distance

### v2

@@ 4.2 Haar wavelet filtering
Algorithm 2
  $l\leftarrow 0$, ${\mathbf{X}}^{(0)}\leftarrow{\mathbf{X}}$ and ${\mathbf{Z}}^{(0)}\leftarrow\mathbf{X}^{\top}$
  $\widehat{\mathbf{W}}_{r}^{(0)}\leftarrow\mathbf{M}_{r}$ and $\widehat{\mathbf{W}}_{c}^{(0)}\leftarrow\mathbf{M}_{c}$
  repeat
   ${\mathbf{X}}^{(l+1)}\leftarrow\Psi({\mathbf{X}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)}))$
   ${\mathbf{Z}}^{(l+1)}\leftarrow\Psi({\mathbf{Z}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)}))$
   $\widehat{\mathbf{X}}_{r}^{(l+1)}\leftarrow\left\{\mathbf{r}_{i}^{(l+1)}=\left({\mathbf{X}}_{i,:}^{(l+1)}\right)^{\top}\middle/\left\lVert{\mathbf{X}}_{i,:}^{(l+1)}\right\rVert_{1}\right\}$
   $\widehat{\mathbf{X}}_{c}^{(l+1)}\leftarrow\left\{\mathbf{c}_{j}^{(l+1)}=\left({\mathbf{Z}}_{j,:}^{(l+1)}\right)^{\top}\middle/\left\lVert{\mathbf{Z}}_{j,:}^{(l+1)}\right\rVert_{1}\right\}$
   $\widehat{\mathbf{W}}_{r}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{r}^{(l+1)};\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)}))$
   $\widehat{\mathbf{W}}_{c}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{c}^{(l+1)};\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)}))$
   $l\leftarrow l+1$
  until convergence
To improve the refinement of trees across iterations, we apply a filtering step at every iteration. Specifically, our filters are constructed from Haar wavelets [Mallat, 1989, 1999, Haar, 1911, Daubechies, 1990]. In [Gavish et al., 2010], it was shown that Haar wavelets can be derived adaptively from trees in a straightforward manner. Here, we propose to build the Haar wavelets from the trees inferred through the iterative process. Viewing each sample (resp. feature) as a signal supported on the feature tree (resp. sample tree) [Ortega et al., 2018] allows us to apply data-driven Haar wavelet filters. Since by construction the filters reflect the intrinsic hierarchical structure of the data, they enhance this meaningful representation while suppressing noise and other nuisance components.
@@
The proof of Thm. 4.2 is in App. C. Same as Thm. 4.1, Thm. 4.2 implies the existence of a limit point to which the proposed iterative scheme with the addition of the Haar wavelet filters converges. We argue that the resulting data-driven wavelet filters attenuate noise and other nuisance components [Gavish et al., 2010], and therefore improve the quality of the learned trees. As shown in Sec. 6, this filtering step contributes to more informative hierarchical representations underlying the high-dimensional data, resulting in superior performance on various tasks. We note that while constructing the filters using $L_{1}$-based selection criterion [Coifman and Wickerhauser, 1992] is effective, alternative filtering strategies [Cohen et al., 1992, Hammond et al., 2011, Shuman et al., 2013, Tremblay and Borgnat, 2014, Leonardi and Van De Ville, 2013, Dong et al., 2016] can be considered, depending on the downstream tasks. We leave this extension for future work.
We conclude this section with a few remarks. First, while our iterative procedure is broadly applicable to various TWD methods [Indyk and Thaper, 2003, Evans and Matsen, 2012, Yamada et al., 2022, Le et al., 2019, Chen et al., 2024], the theoretical results and the ability to obtain meaningful hierarchical representations rely on our choice of the specific tree construction method [Lin et al., 2025]. Second, in each iteration, our method can be computed in $O(n^{1.2}+m^{1.2})$ [Lin et al., 2025, Shen and Wu, 2022] (see App. E). In contrast, a naïve computation requires $O(mn^{3}+m^{3}\log m+nm^{3}+n^{3}\log n)$, making our method more efficient. Third, similar to WSV [Huizing et al., 2022] and Tree-WSV [Düsterwald et al., 2025], our approach can be viewed as an unsupervised ground metric learning technique [Cuturi and Avis, 2014]. However, its specific focus on learning hierarchical representations for both samples and features distinguishes it from these methods and leads to superior empirical performance for hierarchical data (see Sec. 6). For further comparison with Tree-WSV, we refer to App. B and App. G.3. Finally, we remark that either the sample or the feature hierarchical structure can be provided as a prior and used for initialization of our method (see Sec. 5).
@@ 5 Incorporating learned hierarchical representation within HGCNs
In Sec. 4, we introduced a joint hierarchical representation learning framework using TWD. Typically, the initial pairwise distance matrices $\mathbf{M}_{r}$ and $\mathbf{M}_{c}$ are data-driven, e.g., using standard metrics such as the Euclidean distance or the cosine similarity. An advantage of our method is that it can incorporate prior knowledge when a hierarchical structure is available for one of the modes. Without loss of generality, we consider scenarios where a hierarchical structure over the $n$ samples is known and represented by a graph $H=([n],E,\mathbf{A})$, with $[n]=\{1,\ldots,n\}$. To integrate this prior, we initialize $\mathbf{M}_{r}$ using the shortest-path distances $d_{H}$ induced by $H$. This initialization introduces a structured prior in one mode while allowing the hierarchy in the other mode to be learned jointly via TWD.
Incorporating such a prior is particularly relevant for hierarchical graph data [Sen et al., 2008, Chami et al., 2019, Anderson and May, 1991], where the node hierarchy is provided, while the feature structure remains implicit. We demonstrate the compatibility of our approach with hyperbolic graph convolutional networks (HGCNs) [Chami et al., 2019, Dai et al., 2021]. Specifically, during the neighborhood aggregation step in HGCNs, we replace the predefined hierarchical graph with the sample tree inferred by our method after convergence. This learned tree guides the aggregation process, enabling hierarchy-aware message passing that reflects structured relations among samples and across features. As shown in Sec. 6, incorporating our method as a pre-processing step improves performance on link prediction (LP) and node classification (NC) for hierarchical graph datasets. While our focus here is on HGCNs due to their improved performance on NC and LP, the same initialization strategy can be adopted in other hyperbolic architectures, such as hyperbolic neural networks (HNNs) [Ganea et al., 2018, Shimizu et al., 2020]. Further technical details on this integration are provided in App. D.
@@ 6.1 Evaluating hierarchical representations via sparse approximation
We first demonstrate the advantages of the learned trees from Alg. 1 and Alg. 2 for sparse approximation tasks on the data matrix. The quality of the feature tree (and similarly, the sample tree) is evaluated by the $L_{1}$ norm of their expansion coefficients across all samples (and the features, respectively) [Starck et al., 2010, Cohen et al., 2001, Candès et al., 2006]. A lower $L_{1}$ norm indicates a more efficient (sparser) representation of the data using fewer significant Haar coefficients, thus indicating the learned tree structures better reflect the hierarchical information of the data [Gavish et al., 2010, Gavish and Coifman, 2012]. We test four word-document datasets in [Kusner et al., 2015]: BBCSPORT, TWITTER, CLASSIC, and AMAZON, and two scRNA-seq datasets in [Dumitrascu et al., 2021]: ZEISEL and CBMC. Both types of data exhibit hierarchical structures in their features and samples [Blei et al., 2010, Cao et al., 2019]. Additional information about these datasets can be found in App. E.
@@
We propose to learn hierarchical representations for both samples and features, where each informs the other through TWD. To the best of our knowledge, this approach to hierarchical representation learning has not been previously explored. To demonstrate its effectiveness, we compare Alg. 1 and Alg. 2 against existing TWD-based methods as follows. For a fair comparison, we adapt each baseline TWD method to our iterative setting. Specifically, we begin by using the competing method to compute the sample TWD matrix, from which the sample tree is constructed to approximate the corresponding tree-based Wasserstein ground metric. This sample tree is then used to compute the feature TWD matrix, which in turn defines the feature tree. At each iteration, the same baseline method is used to compute the TWDs for samples and features. The competing TWD methods include: Quadtree [Indyk and Thaper, 2003], Flowtree [Backurs et al., 2020], TSWD [Le et al., 2019], UltraTree [Chen et al., 2024], weighted cluster TWD (WCTWD), weighted Quadtree TWD (WQTWD) [Yamada et al., 2022], their sliced variants SWCTWD and SWQTWD, MST [Prim, 1957], Tree Representation (TR) [Sonthalia and Gilbert, 2020], gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) [Monath et al., 2019], gradient-based Ultrametric Fitting (UltraFit) [Chierchia and Perret, 2019], and HC by hyperbolic Dasgupta’s cost (HHC) [Chami et al., 2020]. See App. B for details on these methods. We use the prefix “co-” to denote the adaptation of each method to our iterative framework. In addition, we include comparisons with co-manifold learning that involves trees induced by a diffusion embedding (QUE) [Ankenman, 2014], and Tree-WSV [Düsterwald et al., 2025], which learns unsupervised ground metrics based on tree approximation of WSV [Huizing et al., 2022].
Tab. 1 reports the $L_{1}$ norm of the Haar coefficients across all samples and all features, respectively. We report the value after convergence for our methods, and for baselines, we either report it after convergence or, if convergence is
not achieved, the obtained value after 25 iterations. We see that our methods provide a more efficient representation, showing that the sparsity and quality of the trees produced by our methods are superior and outperform baselines. Fig. 2 shows the $L_{1}$ norm of the Haar expansion coefficients obtained by the proposed methods across iterations on ZEISEL dataset, where we observe that the $L_{1}$ norm is iteratively reduced and reaches convergence. Note that this is not the objective we are minimizing but a consequence of our method of learning well hierarchical representations of the data. Notably, Alg. 2 consistently achieves better sparse approximation performance than Alg. 1. We attribute this improvement to the wavelet filtering step, which jointly considers the data and the structure and further improves the quality of the learned hierarchies at each iteration. One might ask whether wavelet filters could similarly benefit other TWD-based baselines. In App. F, we test this by applying wavelets using trees constructed from various TWD baselines. The results show that it does not consistently improve the quality of the tree representations. We argue that this is because the trees in these methods are primarily designed to approximate the Wasserstein distance as the ground metric, rather than to represent the hierarchical structure of the data. Therefore, these trees are not faithful hierarchical representations of the data, and thus, applying wavelet filters that depend on both data and tree structure fails to improve hierarchical representation learning.
6.2 Document and single-cell classification using learned TWD
We further demonstrate the effectiveness of the TWDs obtained by our methods through document and cell classification tasks. We compare our results with the same competing methods used in Sec. 6.1, and additionally include WSV [Huizing et al., 2022] that learns unsupervised ground metrics as a baseline. Classification is performed using $k$NN based on the obtained distances, with cross-validation over five trials. Each trial randomly splits the dataset into $70\%$ training and $30\%$ testing sets.
@@ 6.2 Document and single-cell classification using learned TWD
Co-Quadtree 96.2$\pm$0.4 69.6$\pm$0.3 95.9$\pm$0.2 89.4$\pm$0.2 81.7$\pm$1.0 80.7$\pm$0.3
Co-Flowtree 95.7$\pm$0.9 71.5$\pm$0.7 95.6$\pm$0.5 91.4$\pm$0.4 84.3$\pm$0.7 83.0$\pm$1.2
Co-WCTWD 93.2$\pm$1.2 70.2$\pm$2.1 94.7$\pm$2.6 87.4$\pm$1.0 82.5$\pm$2.9 79.4$\pm$2.1
Co-WQTWD 95.7$\pm$1.8 70.7$\pm$2.2 95.5$\pm$1.3 88.2$\pm$2.1 82.3$\pm$3.1 80.5$\pm$2.8
Co-UltraTree 95.3$\pm$1.4 70.1$\pm$2.8 93.6$\pm$2.0 86.5$\pm$2.8 85.8$\pm$1.1 84.6$\pm$1.3
Co-TSWD-1 88.2$\pm$1.4 70.4$\pm$1.2 94.7$\pm$0.9 86.1$\pm$0.5 80.2$\pm$1.4 73.2$\pm$1.0
Co-TSWD-5 88.7$\pm$1.7 71.0$\pm$1.5 96.7$\pm$0.8 91.5$\pm$0.4 82.0$\pm$0.9 75.4$\pm$0.7
Co-TSWD-10 89.2$\pm$1.1 71.4$\pm$1.8 95.5$\pm$0.2 91.8$\pm$0.7 83.8$\pm$0.5 77.2$\pm$0.9
Co-SWCTWD 93.5$\pm$2.4 70.5$\pm$1.0 94.4$\pm$1.3 90.7$\pm$1.5 82.7$\pm$1.7 79.0$\pm$0.9
Co-SWQTWD 96.2$\pm$1.2 72.4$\pm$2.1 96.0$\pm$1.1 90.6$\pm$2.3 82.4$\pm$1.4 81.3$\pm$1.1
Co-MST-TWD 88.7$\pm$2.4 68.4$\pm$3.3 91.3$\pm$2.9 87.1$\pm$1.4 80.1$\pm$2.8 76.5$\pm$1.3
Co-TR-TWD 89.5$\pm$1.2 70.9$\pm$1.7 93.4$\pm$2.2 89.5$\pm$1.4 80.7$\pm$0.8 78.5$\pm$0.9
Co-HHC-TWD 86.1$\pm$2.1 70.1$\pm$1.3 93.6$\pm$1.5 88.5$\pm$0.5 83.2$\pm$1.4 77.6$\pm$0.8
Co-gHHC-TWD 84.0$\pm$2.0 70.4$\pm$1.6 90.7$\pm$1.7 87.2$\pm$1.9 79.9$\pm$1.4 84.2$\pm$1.2
Co-UltraFit-TWD 86.8$\pm$0.9 70.9$\pm$1.1 91.9$\pm$1.0 89.9$\pm$2.0 83.7$\pm$2.9 79.1$\pm$1.8
QUE 84.7$\pm$0.5 72.4$\pm$0.6 91.9$\pm$0.5 91.6$\pm$0.9 83.6$\pm$1.4 82.5$\pm$1.9
WSV 85.9$\pm$1.0 71.4$\pm$1.3 92.6$\pm$0.7 89.0$\pm$1.5 81.6$\pm$2.4 77.5$\pm$1.7
Tree-WSV 86.3$\pm$1.5 71.2$\pm$1.9 92.4$\pm$1.0 88.7$\pm$1.9 82.0$\pm$2.9 76.4$\pm$2.4
Alg. 1 96.7$\pm$0.3 74.1$\pm$0.5 97.3$\pm$0.2 94.0$\pm$0.4 90.1$\pm$0.4 86.7$\pm$0.5
Alg. 2 97.3$\pm$0.5 76.7$\pm$0.7 97.6$\pm$0.1 94.2$\pm$0.2 94.0$\pm$0.6 93.3$\pm$0.7

### v3

@@ 4.2 Haar wavelet filtering
To improve the refinement of trees across iterations, we apply a filtering step at every iteration. Specifically, our filters are constructed from Haar wavelets [38, 39, 40, 41]. It was shown that Haar wavelets can be derived adaptively from trees [43]. Here, we propose to build the Haar wavelets from the trees inferred through the iterative process. Viewing each sample (resp. feature) as a signal supported on the feature tree (resp. sample tree) [42] allows us to apply data-driven Haar wavelet filters. Since by construction the filters reflect the intrinsic hierarchical structure of the data, they enhance this meaningful representation while suppressing noise and other nuisance components.
@@
Algorithm 2 Haar Filtering over Alternating Tree Refinement
$l\leftarrow 0$, ${\mathbf{X}}^{(0)}\leftarrow{\mathbf{X}}$, ${\mathbf{Z}}^{(0)}\leftarrow\mathbf{X}^{\top}$, $\widehat{\mathbf{W}}_{r}^{(0)}\leftarrow\mathbf{M}_{r}$ and $\widehat{\mathbf{W}}_{c}^{(0)}\leftarrow\mathbf{M}_{c}$ $\triangleright$ Initialization
repeat
${\mathbf{X}}^{(l+1)}\leftarrow\Psi({\mathbf{X}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)}))$ and ${\mathbf{Z}}^{(l+1)}\leftarrow\Psi({\mathbf{Z}}^{(l)};\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)}))$ $\triangleright$ Tree haar wavelet filtering
$\widehat{\mathbf{X}}_{r}^{(l+1)}\leftarrow\left\{\mathbf{r}_{i}^{(l+1)}=\left({\mathbf{X}}_{i,:}^{(l+1)}\right)^{\top}/\|{\mathbf{X}}_{i,:}^{(l+1)}\|_{1}\right\}$
$\widehat{\mathbf{X}}_{c}^{(l+1)}\leftarrow\left\{\mathbf{c}_{j}^{(l+1)}=\left({\mathbf{Z}}_{j,:}^{(l+1)}\right)^{\top}/\|{\mathbf{Z}}_{j,:}^{(l+1)}\|_{1}\right\}$
$\widehat{\mathbf{W}}_{r}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{r}^{(l+1)};\mathcal{T}(\widehat{\mathbf{W}}_{c}^{(l)}))$ and $\widehat{\mathbf{W}}_{c}^{(l+1)}\leftarrow\Phi(\widehat{\mathbf{X}}_{c}^{(l+1)};\mathcal{T}(\widehat{\mathbf{W}}_{r}^{(l)}))$ $\triangleright$ Iterative update
$l\leftarrow l+1$
until convergence
The proof of Thm. 2 is in App. C. Same as Thm. 1, Thm. 2 implies the existence of a limit point to which the proposed iterative scheme with the addition of the Haar wavelet filters converges. We argue that the resulting data-driven wavelet filters attenuate noise and other nuisance components [43], and therefore improve the quality of the learned trees. As shown in Sec. 6, this filtering step contributes to more informative hierarchical representations underlying the high-dimensional data, resulting in superior performance on various tasks. We note that while constructing the filters using $L_{1}$-based selection criterion [69] is effective, alternative filtering strategies [70, 71, 72, 73, 74, 75] can be considered, depending on the downstream tasks. We leave this extension for future work.
We conclude this section with a few remarks. First, while our iterative procedure is broadly applicable to various TWD methods [32, 33, 34, 35, 36], the theoretical results and the ability to obtain meaningful hierarchical representations rely on our choice of the specific tree construction method [37]. Second, in each iteration, our method can be computed in $O(n^{1.2}+m^{1.2})$ [37, 76] (see App. E). In contrast, a naïve computation requires $O(mn^{3}+m^{3}\log m+nm^{3}+n^{3}\log n)$, making our method more efficient. Third, similar to WSV [59] and Tree-WSV [60], our approach can be viewed as an unsupervised ground metric learning technique [77]. However, its specific focus on learning hierarchical representations for both samples and features distinguishes it from these methods and leads to superior empirical performance for hierarchical data (see Sec. 6). For further comparison with Tree-WSV, we refer to App. B and App. G.3. Finally, we remark that either the sample or the feature hierarchical structure can be provided as a prior and used for initialization of our method (see Sec. 5).
@@ 5 Incorporating learned hierarchical representation within HGCNs
In Sec. 4, we introduced a joint hierarchical representation learning framework using TWD. Typically, the initial pairwise distance matrices $\mathbf{M}_{r}$ and $\mathbf{M}_{c}$ are data-driven, e.g., using standard metrics such as the Euclidean distance or the cosine similarity. An advantage of our method is that it can incorporate prior knowledge when a hierarchical structure is available for one of the modes. Without loss of generality, we consider scenarios where a hierarchical structure over the $n$ samples is known and represented by a graph $H=([n],E,\mathbf{A})$, with $n]=\{1,\ldots,n\}$. To integrate this prior, we initialize $\mathbf{M}_{r}$ using the shortest-path distances $d_{H}$ induced by $H$. This initialization introduces a structured prior in one mode while allowing the hierarchy in the other mode to be learned jointly via TWD.
Incorporating such a prior is particularly relevant for hierarchical graph data [9, 21, 78], where the node hierarchy is provided, while the feature structure remains implicit. We demonstrate the compatibility of our approach with hyperbolic graph convolutional networks (HGCNs) [21, 45]. Specifically, during the neighborhood aggregation step in HGCNs, we replace the predefined hierarchical graph with the sample tree inferred by our method after convergence. This learned tree guides the aggregation process, enabling hierarchy-aware message passing that reflects structured relations among samples and across features. As shown in Sec. 6, incorporating our method as a pre-processing step improves performance on link prediction (LP) and node classification (NC) for hierarchical graph datasets. While our focus here is on HGCNs due to their improved performance on NC and LP, the same initialization strategy can be adopted in other hyperbolic architectures, such as hyperbolic neural networks (HNNs) [79, 80]. Further technical details on this integration are provided in App. D.
@@ 6.1 Evaluating hierarchical representations via sparse approximation
We first demonstrate the advantages of the learned trees from Alg. 1 and Alg. 2 for sparse approximation tasks on the data matrix. The quality of the feature tree (and similarly, the sample tree) is evaluated by the $L_{1}$ norm of their expansion coefficients across all samples (and the features, respectively) [81, 82, 83]. A lower $L_{1}$ norm indicates a more efficient (sparser) representation of the data using fewer significant Haar coefficients, thus indicating the learned tree structures better reflect the hierarchical information of the data [43, 53]. We test four word-document datasets [25]: BBCSPORT, TWITTER, CLASSIC, and AMAZON, and two scRNA-seq datasets [84]: ZEISEL and CBMC. Both types of data exhibit hierarchical structures in their features and samples [85, 86]. In document data, words (features) form semantic hierarchies (e.g., animal → mammal → dog), while documents (samples) follow topic–subtopic structures (e.g., science → biology → genomics). In scRNA-seq data, genes (features) are organized by functional relationships such as biological pathways and gene ontologies (e.g., immune response genes → cytokine genes → specific interleukins), and cells (samples) follow developmental or taxonomic hierarchies (e.g., hematopoietic stem cell → progenitor cell → mature blood cell types), which are widely modeled as hierarchical in computational biology [87]. Additional information about these datasets can be found in App. E.
@@
[…]
We propose to learn hierarchical representations for both samples and features, where each informs the other through TWD. To the best of our knowledge, this approach to hierarchical representation learning has not been previously explored. To demonstrate its effectiveness, we compare Alg. 1 and Alg. 2 against existing TWD-based methods as follows. For a fair comparison, we adapt each baseline TWD method to our iterative setting. Specifically, we begin by using the competing method to compute the sample TWD matrix, from which the sample tree is constructed to approximate the corresponding tree-based Wasserstein ground metric. This sample tree is then used to compute the feature TWD matrix, which in turn defines the feature tree. At each iteration, the same baseline method is used to compute the TWDs for samples and features. The competing TWD methods include: Quadtree [32], Flowtree [88], TSWD [35], UltraTree [36], weighted cluster TWD (WCTWD), weighted Quadtree TWD (WQTWD) [34], their sliced variants SWCTWD and SWQTWD, MST [89], Tree Representation (TR) [22], gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) [90], gradient-based Ultrametric Fitting (UltraFit) [91], and HC by hyperbolic Dasgupta’s cost (HHC) [92]. See App. B for details on these methods. We use the prefix “co-” to denote the adaptation of each method to our iterative framework. In addition, we include comparisons with co-manifold learning that involves trees induced by a diffusion embedding (QUE) [52], and Tree-WSV [60], which learns unsupervised ground metrics based on tree approximation of WSV [59].
Tab. 1 reports the $L_{1}$ norm of the Haar coefficients across all samples and all features, respectively. We report the value after convergence for our methods, and for baselines, we either report it after convergence or, if convergence is not achieved, the obtained value after 25 iterations. We see that our methods provide a more efficient representation, showing that the sparsity and quality of the trees produced by our methods are superior and outperform baselines. Fig. 2 shows the $L_{1}$ norm of the Haar expansion coefficients obtained by the proposed methods across iterations on ZEISEL dataset, where we observe that the $L_{1}$ norm is iteratively reduced and reaches convergence. Note that this is not the objective we are minimizing but a consequence of our method of learning well hierarchical representations of the data. Notably, Alg. 2 consistently achieves better sparse approximation performance than Alg. 1. We attribute this improvement to the wavelet filtering step, which jointly considers the data and the structure and further improves the quality of the learned hierarchies at each iteration. One might ask whether wavelet filters could similarly benefit other TWD-based baselines. In App. F, we test this by applying wavelets using trees constructed from various TWD baselines. The results show that it does not consistently improve the quality of the tree representations. We argue that this is because the trees in these methods are primarily designed to approximate the Wasserstein distance as the ground metric, rather than to represent the hierarchical structure of the data. Therefore, these trees are not faithful hierarchical representations of the data, and thus, applying wavelet filters that depend on both data and tree structure fails to improve hierarchical representation learning.
6.2 Document and single-cell classification using learned TWD
We further demonstrate the effectiveness of the TWDs obtained by our methods through document and cell classification tasks. We compare our results with the same competing methods used in Sec. 6.1, and additionally include WSV [59] that learns unsupervised ground metrics as a baseline. Classification is performed using kkNN based on the obtained distances, with cross-validation over five trials. Each trial randomly splits the dataset into $70\%$ training and $30\%$ testing sets.